# Granger Causality in High-Dimensional Networks of Time Series

### v2

@@ 4. Fit Restricted/Unrestricted Models:
    Unrestricted Model:
    Restricted Model:
    $Y_{t}=\gamma_{0}+\sum_{i=1}^{p}\gamma_{i}Y_{t-i}+\eta_{t}$
    5. Perform Granger Causality Tests:
@@
      Test null hypothesis $H_{0}:\beta_{j}=0$
    for all $j$ in the equation for $Y_{t}$.
  • ii
    Compute F-statistic:
    $F=\frac{(RSS_{R}-RSS_{U})/q}{RSS_{U}/(T-K)}$
    where:
    • $RSS_{R}$: Residual sum of squares of
  • restricted model (without $X_{t-j}$ terms).
  • $RSS_{U}$: Residual sum of squares of
  • unrestricted model.
  • $q$: Number of restrictions (parameters set to zero). $K$: Total number of parameters in the unrestricted model. iii Compare F-statistic with critical value
  • from $F$-distribution $F_{q,T-K}$.
  • iv
    Reject $H_{0}$ if $F$ exceeds critical value; conclude $X_{t}$ Granger-causes $Y_{t}$. v Repeat steps for testing causality from $Y_{t}$ to $X_{t}$.
@@
    III Simulation study
    The aim of this simulation study is to evaluate the effectiveness of spectral dynamic principal component analysis (sDPCA) combined with Granger causality (GC) testing in identifying true causal relationships within a high-dimensional network of time series data, specifically focusing on EEG-like data. The study investigates how well the proposed method isolates the channels of interest (COI) from the influences of other channels and accurately assesses their causal interactions. We conduct two sets of simulations: one with linear influences and another with nonlinear influences.
    III-A Linear Scheme
    In the linear scheme of the simulation study, we generate a network of time series data with predefined causal relationships and linear influences from additional channels.
    III-A1 Generating Initial Channels
    We begin by generating $N=40$ initial time series channels, denoted by $\{Z_{i}(t)\}_{i=1}^{N}$, over time points $t=1,\ldots,T$. The channels are divided into $N/2=20$ causal pairs, where each pair consists of a ”cause” and an ”effect” channel.
    For each cause channel $Z_{i}(t)$, where $i=1,\ldots,N/2$, we model it as a weighted sum of latent processes representing different EEG frequency bands plus noise:
    $Z_{i}(t)=a_{i}\,\delta(t)+b_{i}\,\theta(t)+c_{i}\,\alpha(t)+d_{i}\,\beta(t)+e_{i}\,\gamma(t)+\varepsilon_{i}(t),$
    where:
    • $\delta(t)$, $\theta(t)$, $\alpha(t)$, $\beta(t)$, and $\gamma(t)$ are latent processes corresponding to the EEG frequency bands:
      • Delta (0.50.50.50.54444 Hz): $\delta(t)$,
      • Theta (44448888 Hz): $\theta(t)$,
      • Alpha (888812121212 Hz): $\alpha(t)$,
      • Beta (1212121230303030 Hz): $\beta(t)$,
      • Gamma ($>$ $30$ Hz): $\gamma(t)$.
      Each latent process is modeled as an autoregressive process of order 2 (AR(2)):
      $\nu(t)=\phi_{1\nu}\nu(t-1)+\phi_{2\nu}\nu(t-2)+\eta_{\nu}(t),\quad\nu\in\{\delta,\theta,\alpha,\beta,\gamma\},$
      where $\eta_{\nu}(t)$ is white noise. The AR coefficients can expressed in terms of $r_{\nu}$ and $\varrho_{\nu}$:
      $\phi_{1\nu}=2r_{\nu}\cos(\varrho_{\nu}),$
      $\phi_{2\nu}=-r_{\nu}^{2}.$
      where $r_{\nu}$ is the damping factor, chosen close to $1$ (e.g., $10.8\leq r_{\nu}<$ 1) to ensure stationarity and control the bandwidth of the spectral peak. And, $\varrho_{\nu}$ is the angular frequency (in radians per sample), related to the desired frequency $f_{\nu}$ (in Hz) by:
      $\varrho_{\nu}=2\pi\frac{f_{\nu}}{f_{s}},$
      with $f_{s}$ being the sampling frequency. By modeling the latent processes to correspond to specific EEG frequency bands, we aim to simulate EEG-like signals more realistically.
    • $a_{i}$, $b_{i}$, $c_{i}$, $d_{i}$, and $e_{i}$ are random weights drawn from a uniform distribution to ensure variability among channels.
    • $\varepsilon_{i}(t)$ is white noise added to each channel.
    For each effect channel $Z_{j}(t)$, where $j=i+N/2$ and $i=1,\ldots,N/2$, we generate it as an autoregressive process influenced by its corresponding cause channel:
    $Z_{j}(t)=\sum_{k=1}^{p}\phi_{k}Z_{i}(t-k)+\varepsilon_{j}(t),$
    where:
    • $p=2$ is the order of the autoregressive process.
    • $\phi_{k}$ are AR coefficients ensuring the stability of the process.
    • $\varepsilon_{j}(t)$ is white noise added to the effect channel.
    III-A2 Extending the Network
    We extend the network by generating additional $N_{\text{ext}}=472$ channels, resulting in a total of $N_{\text{expanded}}=512$ channels. These additional channels $\{Z_{i}(t)\}_{i=N+1}^{N_{\text{expanded}}}$ are generated similarly to the cause channels but without being part of the predefined causal pairs:
    $Z_{i}(t)=a_{i}\,\delta(t)+b_{i}\,\theta(t)+c_{i}\,\alpha(t)+d_{i}\,\beta(t)+e_{i}\,\gamma(t)+\varepsilon_{i}(t).$
    III-A3 Introducing Linear Influences
    To simulate the effect of a high-dimensional network on the initial channels, we introduce linear influences from a subset of the additional channels. We select a set of connected nodes $\mathcal{C}\subset\{N+1,\ldots,N_{\text{expanded}}\}$ of size nofcn (number of connected nodes).
@@ III-A3 Introducing Linear Influences
    For each initial channel $Z_{i}(t)$, $i=1,\ldots,N$, we randomly select $k$ influencing nodes from $\mathcal{C}$ and assign influence weights $W_{ij}$ drawn from a uniform distribution.
    The linear influence on channel $i$ at time $t$ is defined as:
    $L_{i}(t)=\sum_{j\in\mathcal{I}_{i}}W_{ij}Z_{j}(t),$
    where $\mathcal{I}_{i}$ is the set of influencing nodes for channel $i$. For each initial channel $Z_{i}(t)$ and each influencing node $Z_{j}(t)\in\mathcal{I}_{i}$, we assign an influence weight $W_{ij}$ drawn from a uniform distribution centered around a specified influence weight $\omega$:
    $W_{ij}\sim\text{Uniform}\left(\omega-\frac{\omega}{16},\,\omega+\frac{\omega}{16}\right),\quad\text{for }j\in\mathcal{I}_{i}.$
    We then update the initial channels by adding the linear influence:
    $\tilde{Z}_{i}(t)=Z_{i}(t)+L_{i}(t).$
    The updated initial channels $\tilde{Z}_{i}(t)$ now contain both their original dynamics and the linear influence from the connected nodes.
    III-B Nonlinear Scheme
    In the nonlinear scheme, we introduce moderate nonlinearity into the influence of the connected nodes on the initial channels to study the robustness of the sDPCA and GC testing framework under nonlinear conditions.
    III-B1 Introducing Nonlinear Influences
    Similar to the linear simulation, we select a set of connected nodes $\mathcal{C}$ and assign influence weights $W_{ij}$ for each initial channel $Z_{i}(t)$.
    The linear influence is computed as before:
    $L_{i}(t)=\sum_{j\in\mathcal{I}_{i}}W_{ij}Z_{j}(t).$
    To introduce nonlinearity, we compute interaction terms between the influencing nodes for each initial channel:
    $Q_{i}(t)=\sum_{j\in\mathcal{I}_{i}}\sum_{k>j}W_{ij}W_{ik}Z_{j}(t)Z_{k}(t).$
    We apply a nonlinear function $f(\cdot)$ to the linear influence, such as the softplus function:
    $f(x)=\ln(1+e^{x}).$
    The total influence on channel $i$ at time $t$ is then given by:
    $S_{i}(t)=f\left(L_{i}(t)\right)+\omega Q_{i}(t),$
    where $\omega$ is a scaling parameter controlling the magnitude of the nonlinear interaction influence.
    We update the initial channels with the total influence:
    $\tilde{Z}_{i}(t)=Z_{i}(t)+S_{i}(t).$
    III-C Removing Influences
    III-C1 Removing Influences Using sDPCA
    To isolate the channels of interest from the influences of the connected nodes, we apply sDPCA to the time series data of the connected nodes $\{Z_{j}(t)\}_{j\in\mathcal{C}}$. We obtain principal component scores $\{Z_{\text{pc},k}(t)\}_{k=1}^{N_{\text{dpc}}}$, where $N_{\text{dpc}}$ is the number of dynamic principal components retained. For each adjusted initial channel $\tilde{Z}_{i}(t)$, we regress it on the principal component scores and their interaction terms:
    $\tilde{Z}_{i}(t)=\sum_{k=1}^{N_{\text{dpc}}}\varphi_{ik}Z_{\text{pc},k}(t)+\sum_{k\neq l}\psi_{ikl}Z_{\text{pc},k}(t)Z_{\text{pc},l}(t)+\epsilon_{i}(t).$
    We then use the residuals $\hat{\epsilon}_{i}(t)$ as the adjusted time series for further analysis:
    $\hat{Z}_{i}(t)=\hat{\epsilon}_{i}(t).$
    III-C2 Granger Causality Testing
    Using the adjusted time series $\{\hat{Z}_{i}(t)\}_{i=1}^{N}$, we fit vector autoregressive (VAR) models and perform Granger causality tests to identify causal relationships among the initial channels. The effectiveness of the sDPCA adjustment is evaluated by comparing the detected causal relationships with the ground truth.
    III-D Simulation Results
    To evaluate the performance of the proposed method, we compute accuracy metrics based on the detection of true causal relationships. Specifically, we compare the identified causal links with the known ground truth and compute measures such as accuracy rate.
    By conducting both linear and nonlinear simulations, we assess the robustness of the sDPCA combined with Granger causality testing in high-dimensional settings with complex interdependencies. The results provide insights into the method’s effectiveness in isolating the channels of interest and accurately detecting causal relationships, even in the presence of nonlinear influences.
    The results are sensitive to the choice of latent processes and the influence of other channels. Different scenarios are explored by varying the number of influencing channels and the weight of their influence, with the goal of assessing the robustness and reliability of the proposed method in detecting true causal relationships within high-dimensional time series networks. For example, Table I presents the accuracy results for one of the simulation studies for different numbers of influencing nodes and varying weights on the COI in a high-dimensional network. The COI consists of the first 40 nodes out of a total of 512 nodes. The results of the other 5 simulation studies are given in the appendix.
    TABLE I: Accuracy results for varying numbers of influencing nodes with nonlinear influence and weights on the COI in a high-dimensional network under a non-normal error distribution, $\epsilon\sim\mathrm{non\text{-}Norm}(\mu=-1,\sigma=1,\gamma=1.5,\kappa=3.5).$ The COI consists of the first 40 nodes out of a total of 512 nodes. The numbers 30, 60, 90, and 120 represent the number of nodes influencing the COI, and the values 0.1, 0.3, 0.5, and 0.7 represent the weights of this influence.
    Influencing Nodes = 30
    Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
    3 0.96 0.88 0.61 0.65
    6 0.97 0.77 0.62 0.64
    12 0.97 0.82 0.61 0.65
    24 0.93 0.82 0.65 0.64
    Influencing Nodes = 60
    Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
    3 0.98 0.66 0.64 0.64
    6 0.99 0.73 0.64 0.65
    12 0.94 0.69 0.64 0.64
    24 0.96 0.67 0.64 0.64
    Influencing Nodes = 90
    Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
    3 0.96 0.63 0.63 0.65
    6 0.94 0.62 0.64 0.64
    12 0.95 0.62 0.64 0.64
    24 0.91 0.64 0.64 0.64
    Influencing Nodes = 120
    Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
    3 0.95 0.65 0.63 0.64
    6 0.95 0.65 0.64 0.65
    12 0.90 0.64 0.65 0.64
    24 0.97 0.64 0.64 0.64
    This simulation evaluates the performance of the sDPCA method under a nonlinear influence scheme with non-normal error terms, introducing complexity due to both nonlinearity and deviations from normality.
    At a low influence weight (weight = 0.1), sDPCA maintains high accuracy across all numbers of influencing nodes, even as the network becomes super connected and super dense. Accuracy exceeds 0.90 when the number of influencing nodes increases from 30 up to 120, indicating strong robustness in isolating the channels of interest (COI) even in highly interconnected networks where each node affects the causal pairs.
    As the influence weight increases to 0.3, accuracy decreases but remains relatively high with fewer influencing nodes. For example, with 30 influencing nodes and Ndpc = 3, the accuracy is 0.88. However, as the network becomes denser (influencing nodes ¿ 30), the decline in accuracy becomes more pronounced, reflecting the increased challenge of isolating true causal relationships in a super connected network.
    At higher influence weights (weight = 0.5 and 0.7), accuracy declines further and stabilizes around 0.64–0.65, regardless of the number of influencing nodes or the number of dynamic principal components (Ndpc). This plateau suggests a limit to the method’s ability to isolate the COI under strong nonlinear influences and pervasive connectivity, as heavier weights and a denser network make it more challenging to distinguish true causal relationships from spurious ones introduced by nonlinearity and widespread influence.
    The impact of increasing Ndpc on accuracy is minimal at higher weights and in denser networks, indicating that adding more principal components does not enhance performance under these conditions. This suggests that the sDPCA method’s capacity to capture and remove complex influences is limited when the network is both highly connected and subject to strong nonlinear interactions.
    Overall, the sDPCA method demonstrates robustness to nonlinear influences and non-normal error distributions at lower influence weights and in less densely connected networks. Its effectiveness diminishes as the strength of the nonlinear influence increases and as the network becomes super connected, highlighting the importance of considering both the magnitude of external influences and network connectivity when applying sDPCA in practical settings with nonlinear dynamics.
    IV Application to EEG data
    In this section, we apply our proposed approach to real EEG data from an experiment involving motor execution and imagery tasks, offering insights into the temporal dynamics of brain activity. EEG recordings were collected from 109 volunteers using the BCI2000 system with a 64-channel setup; the data is publicly available at PhysioNet.org [18]. Each subject completed 12 task-oriented runs. Participants either physically performed or imagined movements in response to visual cues on a screen, such as opening and closing fists or feet depending on the target’s position. The tasks were divided into four categories: physical movements of the left or right fist, imaginary movements of the left or right fist, physical movements of both fists or both feet, and imaginary movements of both fists or both feet. EEG data were sampled at 160 Hz, along with an annotation channel indicating rest or the onset of motion (real or imagined) for specific movements. The EEG montage followed the international 10-10 system, omitting certain electrodes (Nz, F9, F10, FT9, FT10, A1, A2, TP9, TP10, P9, and P10).
    All data underwent pre-processing to filter out artifacts like eye blinks and muscle movements, preserving the temporal structure of the recordings. The procedure included a high-pass filter at 0.5 Hz to eliminate low-frequency noise, removal of 60 Hz AC line noise, and Independent Component Analysis (ICA) using MNE-Python routines [19, 20] to isolate and exclude artifact components. Finally, artifacts associated with eye blinking and muscle movements were identified and removed, ensuring the recordings were free from these specific interferences.
    Despite standard pre-processing, some recordings had unresolved issues like sawtooth traces, excessive blinking, broken channels, and unbalanced epochs. Therefore, we limited our analysis to 35 right-handed subjects (17 females and 18 males) whose recordings were qualified. To demonstrate how the proposed method can dissect causal connections between pairwise nodes in any time series network, we focused on six channels of interest (COI): AF3 and AF4 positioned over the pre-frontal cortex, C3 and C4 placed over the motor cortex, and O1 and O2 over the occipital cortex, as illustrated in Figure 1. These channels were chosen because they correspond to brain regions likely to be active during voluntary movement, imagery, and visual perception.

### v3

@@ 4. Fit Restricted/Unrestricted Models:
    $Y_{t}=\alpha_{0}+\sum_{i=1}^{p}\alpha_{i}Y_{t-i}+\sum_{j=1}^{q}\beta_{j}X_{t-j}+\epsilon_{t}$.
  2. ii
    $Y_{t}=\gamma_{0}+\sum_{i=1}^{p}\gamma_{i}Y_{t-i}+\eta_{t}$.
5. Testing Granger Causality:
  1. i
    Test $H_{0}:\beta_{j}=0$ for all $j$.
  2. ii
    Compute F-statistic $F=\frac{(RSS_{R}-RSS_{U})/q}{RSS_{U}/(T-K)}$.
  3. iii
    Reject $H_{0}$ if $F$ exceeds the critical value $F_{q,T-K}$.
@@ 5. Testing Granger Causality:
3 Simulation study
This section presents a framework for evaluating whether Granger causality analysis, combined with sDPCA-based removal of confounding effects, can recover underlying causal associations in a set of EEG-like time series networks. The main objective is to demonstrate how this approach addresses linear, nonlinear, and causative external influences in high-dimensional settings. The third scenario is evaluated in a comparative framework. In this context, GC analysis with conventional PCA, DFM-based approaches by Barigozzi et al. (2024), and sparse vector autoregressive (VAR) estimation through parameter penalization techniques by (Nicholson et al., 2017, 2020) can be summoned up to obtain GC associations in large networks. These methods are made available in the R environment. Although the DFM estimation provided by the fnets package and the penalized VAR estimations provided by the BigVAR package are not directly designed for GC analysis, they can be regarded as alternatives since the estimated sparse coefficient matrices can signify GC links among variables. We have also considered alternatives where tests for GC detection can be performed with these techniques. All simulations incorporate designed connections among specific channels, while outer nodes of networks introduce confounding influences to obscure them.
Suppose the total size of the generated networks is $N+N_{external}$ in which signals are labeled as $X_{1},\ldots,X_{N/2},Y_{1},\ldots,Y_{N/2},Z_{1},\ldots,Z_{N_{\mathrm{ext}}}$. The goal is to explore how well one can detect the causal connection $X_{i}\to Y_{i}$ after these pairs are affected by additional channels $\{Z_{j}\}$. Each $X^{*}_{i}$ is produced by mixing latent oscillatory components that reflect distinct frequency bands commonly used to approximate EEG activity. These bands—delta (0.50.50.50.54444 Hz), theta (44448888 Hz), alpha (888812121212 Hz), beta (1212121230303030 Hz), and gamma ($>$ $30$ Hz)—are treated as AR(2) processes of the form $\nu(t)=\psi_{1\nu}\nu(t-1)+\psi_{2\nu}\nu(t-2)+\eta_{\nu}(t),$ where $\psi_{1\nu}=2r_{\nu}\cos(\varrho_{\nu})$, $\psi_{2\nu}=-r_{\nu}^{2}$, and $\eta_{\nu}(t)\sim\mathrm{WN}(0,\sigma^{2}_{\nu})$. The resulting latent components, for $\nu\in\{\delta,\theta,\alpha,\beta,\gamma\}$, allow each $\nu$ to maintain a stationary and nearly resonant character. A typical $X_{i}$ can be written as
$X_{i}(t)=a_{i}\delta(t)+b_{i}\theta(t)+c_{i}\alpha(t)+d_{i}\beta(t)+e_{i}\gamma(t)+\varepsilon_{i}(t),$
where $\{a_{i},b_{i},c_{i},d_{i},e_{i}\}$ are channel-specific weights and $\varepsilon_{i}\sim\mathrm{WN}(0,\sigma^{2}_{\varepsilon})$. Its paired $Y^{*}_{i}$ is constructed by applying an AR model to $X^{*}_{i}$:
$Y_{i}(t)=\sum_{k=1}^{p}\psi_{k}X_{i}(t-k)+\xi_{i}(t),\text{ where }\xi_{i}(t)\sim\mathrm{WN}(0,\sigma^{2}_{\xi})$
with $p=2$ and stationarity-providing coefficients $\psi_{k}\in\pm 1.5$ in this study. This pairing establishes underlying cause-and-effect association $X_{i}\to Y_{i}$. An extended network is formed by adding exterior channels $\{Z_{j}\}$ to the initial causal setting, each generated in a similar fashion to $X_{i}$ but not designed to have cause-effect connections with the original pairs. These outer signals form a large repository of possible confounders. The final set contains $N_{\text{expanded}}$ time series, with the first $N$ holding the $N/2$ cause-effect relationships and the additional $N_{\text{external}}$ acting as external influencing signals that can inject disturbances into the original structure.
3.1 Linear Scheme
@@
In order to simulate the confounding effect of a high-dimensional network on the initial channels, linear influences are introduced from a subset of the additional channels $\{Z_{j}\}$ to blur the association between $X_{i}$ and $Y_{i}$. We select a set of connected nodes $\mathcal{I}\subset\{1,\ldots,N_{ext}\}$. For each initial channel $Z_{j}(t)$, $j=1,\ldots,N_{ext}$, we randomly select $k$ influencing nodes from $\mathcal{I}$ and assign influence weights $W_{ij}$ drawn from a uniform distribution. Then, the linear influence (i.e., $L_{i}(t)=\sum_{j\in\mathcal{I}_{i}}W_{ij}Z_{j}(t)$) on channel $i$ at time $t$ can be given as:
${X}^{*}_{i}(t)=X_{i}(t)+\sum_{j\in\mathcal{I}_{i}}W_{ij}Z_{j}(t),$
${Y}^{*}_{i}(t)=Y_{i}(t)+\sum_{j\in\mathcal{I}_{i}}W_{ij}Z_{j}(t),$
where $\mathcal{I}_{i}$ is the set of influencing nodes for channel $i$. For each initial channel $X^{*}_{i}(t)$ and each influencing node $Z_{j}(t)\in\mathcal{I}_{i}$, we assign an influence weight $W_{ij}$ drawn from a uniform distribution centered around a specified influence weight $\omega$ ( i.e., $W_{ij}\sim\text{Uniform}(\omega-\frac{\omega}{16},\omega+\frac{\omega}{16})$). The modified signals ${X}^{*}_{i}$ and ${Y}^{*}_{i}$ retain their causal arrangement in principle, but it is obscured by superimposed signals from the outer set. In the case of large amplitude weights (i.e., higher connectivity), the underlying causal association will be difficult to detect unless these background effects are adequately addressed.
3.2 Nonlinear Scheme
In the nonlinear scheme, we introduce moderate nonlinearity into the initial channels. The outer channels $\{Z_{j}\}$ remain potential confounders, but each channel ${X}^{*}_{i}$ or ${Y}^{*}_{i}$ includes products $Z_{j}(t)Z_{k}(t)$ and a smooth function of linear influence $L_{i}(t)=\sum_{j\in\mathcal{I}_{i}}W_{ij}Z_{j}(t)$. The nonlinear influence on channel $i$ at time $t$ can be exemplified as:
${X}^{*}_{i}(t)=X_{i}(t)+f\Bigl{(}\sum_{j\in\mathcal{I}_{i}}W_{ij}Z_{j}(t)\Bigr{)}+\omega^{*}\sum_{j\in\mathcal{I}_{i}}\sum_{k>j}W_{ij}W_{ik}Z_{j}(t)Z_{k}(t),$
${Y}^{*}_{i}(t)=Y_{i}(t)+f\Bigl{(}\sum_{j\in\mathcal{I}_{i}}W_{ij}Z_{j}(t)\Bigr{)}+\omega^{*}\sum_{j\in\mathcal{I}_{i}}\sum_{k>j}W_{ij}W_{ik}Z_{j}(t)Z_{k}(t),$
where $f(x)=\ln\bigl{(}1+e^{x}\bigr{)}$ and $\omega^{*}$ scales the interaction term.
3.3 Causal Scheme
A separate simulation setup imposes cause-effect links among initial nodes and uses a broader set of $\{Z_{j}\}$ to create a complex background. The outer channels modify each $X_{i}$ or $Y_{i}$ according to difference equations like
${X}^{*}_{i}(t)=\mu X_{i}(t-1)+\sum_{j\in\mathcal{I}_{i}}W_{ij}Z_{j}(t-1)+\nu_{i}(t),\text{ where }\nu_{i}(t)\sim\mathrm{WN}(0,\sigma^{2}_{\nu})$
${Y}^{*}_{i}(t)=\mu Y_{i}(t-1)+\sum_{j\in\mathcal{I}_{i}}W_{ij}Z_{j}(t-1)+\nu_{i}(t),\text{ where }\nu_{i}(t)\sim\mathrm{WN}(0,\sigma^{2}_{\nu})$
making it difficult to recover the direct link $X_{i}\to Y_{i}$ unless one accounts for external influences. This form is used as a framework in which multiple strategies can be compared.
3.4 Simulation Results
Linear and Nonlinear Schemes of Influences
We assessed the proposed approach’s ability to recover known causal connections in simulated high-dimensional networks having both linear and nonlinear interference from outer channels explained in 3.1 and 3.2. In these settings, different numbers of influencing nodes (with varying weights) on the COI in a high-dimensional network are introduced. The size of the networks is set to 512 nodes. Thus, the COI consists of the first 40 (i.e., $N=40$) nodes out of a total of 512 (i.e., $N_{expanded}=512$) nodes. Figure 1 summarizes the results across these multiple setups. Results show that the procedure generally maintains high accuracy when confounding is modest. As the influence weight and network density grow, the method’s performance declines and stabilizes, indicating that strong external perturbations and extensive connectivity cause significant challenges. Increasing the number of dynamic principal components beyond a small set does not substantially improve accuracy under heavier confounding. However, the number of scores can be chosen according to their ability for the total variance explainability. In these scenarios, we focused more on a relatively small number of scores rather than explainability, so, total variance explainability changes between $20\%$ to $65\%$. Overall, the results suggest that spectral dynamic principal component analysis can effectively isolate the channels of interest in large networks, though very high connectivity or influence weights reduce its success.
Figure 1: Accuracy results under Linear and Nonlinear Influences
The method achieved high accuracy (above 0.90) when influence weights were low (e.g., 0.1), even as the number of interfering nodes increased from 30 to 120. As the influence weight rose to 0.3, performance remained acceptable for smaller interfering sets but declined with increasing network density.
When the influence weights reached 0.5 and 0.7, the accuracy obtained was around 0.64–0.65, unaffected by the number of dynamic principal components used. This suggests that under strong nonlinear disturbances and widespread or super connectivity, the method’s ability to disentangle the channels of interest becomes limited.
In lower interference regimes, sDPCA consistently filtered out background effects and preserved meaningful causal relationships. However, in highly entangled networks with amplified nonlinear effects, performance deteriorates regardless of the dimensionality reduction depth. These findings point to the method’s strengths in moderately complex environments and its limitations in ultra-dense or strongly nonlinear settings.
Causative Schemes of Influences
In another simulation scenario, we consider the case where the outer network acts in a causative manner given in 3.3. The size of the networks is set to 100 nodes in this case. The COI consists of the first 20 (i.e., $N=20$) nodes out of a total of 100 (i.e., $N_{expanded}=100$) nodes. Figure 2 shows boxplots of the Matthews correlation coefficient (MCC) and Cohen’s Kappa in the upper panels, along with line plots of mean values in the lower panels. Each plot illustrates how well different methods detect the injected causal links under varying degrees of network density (low, moderate, and high connectivity). The horizontal axis indicates the number of influencing nodes, which grows from low to high levels of confounding. GC–PCA uses standard principal component analysis to remove shared background signals before testing for Granger causality. DFM1, the default procedure, extracts idiosyncratic components via dynamic factor modeling and fits a sparse VAR to interpret nonzero coefficients as causal links. DFM2 modifies this by applying a formal Granger test to the idiosyncratic signals, rather than relying solely on a coefficient matrix. BigVAR1 fits an elastic-net penalized VAR to the influenced network, while BigVAR2 uses the same penalization strategy but only on factor-based idiosyncratic components extracted, forming a hybrid of DFM and penalized VAR estimation.
Figure 2: Comparison of accuracies under Causative Influences
Granger causality combined with sDPCA tends to provide higher accuracy scores across all scenarios. GC-PCA also performs well, though becomes less noticeable when the network density increases. Unlike the previous scenario, here, the number of scores used in partialling out of the effect of the outer network was determined when the cumulative explained variance reached $75\%$. Fitting BigVAR to only the idiosyncratic components estimated via DFM (BigVAR2) improves performance relative to fitting it to the full network (BigVAR1). DFM1 (default mode) shows relatively low performance in these simulations, whereas DFM2 (a conventional test for GC on the idiosyncratic components) recovers more causal links but remains below the top-performing procedures. Accuracy measures decrease as the number of influencing nodes grows, reflecting the increased difficulty of recovering true causal connections in super-connected systems. Methods that factor out or model the common background, such as sDPCA or the hybrid approach in BigVAR2, tend to be more robust in this regard.
4 Application to EEG data
In this section, we apply our proposed approach to EEG recordings from an experiment involving motor execution and imagery tasks, offering insights into the temporal dynamics of brain activity. EEG signals were collected from multiple subjects using the BCI2000 system with a 64-channel setup; the data is publicly available at PhysioNet.org (Schalk et al., 2004). Each subject completed 12 task-oriented runs. Participants either physically performed or imagined movements in response to visual cues on a screen, such as opening and closing fists or feet, depending on the target’s position. The tasks were divided into four categories: physical movements of the left or right fist, imaginary movements of the left or right fist, physical movements of both fists or both feet, and imaginary movements of both fists or both feet. EEG data were sampled at 160 Hz, along with an annotation channel indicating rest or the onset of motion (real or imagined) for specific movements. The EEG montage followed the international 10-10 system, omitting certain electrodes (Nz, F9, F10, FT9, FT10, A1, A2, TP9, TP10, P9, and P10).
All data underwent pre-processing to filter out artifacts like eye blinks and muscle movements, preserving the temporal structure of the recordings. The procedure included a high-pass filter at 0.5 Hz to eliminate low-frequency noise, removal of 60 Hz AC line noise, and Independent Component Analysis (ICA) using MNE-Python routines (Gramfort et al., 2013; Larson et al., 2024) to isolate and exclude artifact components.
Despite standard pre-processing, some recordings had unresolved issues like sawtooth traces, excessive blinking, broken channels, and unbalanced epochs. Therefore, we limited our analysis to 35 right-handed subjects (17 females and 18 males) whose recordings were qualified. To demonstrate how the proposed method can dissect causal connections between pairwise nodes in a network, we focused on six channels of interest (COI) out of 64 EEG channels: AF3 and AF4 positioned over the pre-frontal cortex, C3 and C4 placed over the motor cortex, and O1 and O2 over the occipital cortex, as illustrated in Figure 3(a). These channels were chosen because they correspond to brain regions likely to be active during voluntary movement, imagery, and visual perception.